OpenStudy (anonymous):

What is the domain of the square root of the square root of x-1-1

ganeshie8 (ganeshie8):

\(\large \mathbb{ \sqrt{\sqrt{x-1} - 1} }\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

the stuff inside sqrt has to be >= 0

ganeshie8 (ganeshie8):

so, \(\large \mathbb{ \sqrt{x-1} - 1 \ge 0 }\)

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

x is \[x \ge 2\]

ganeshie8 (ganeshie8):

yes ! thats the domain

ganeshie8 (ganeshie8):

^ur answer matches wid wolfram

OpenStudy (anonymous):

does the zero thing usually work with equations like this?

ganeshie8 (ganeshie8):

yes, the assumption is that all functions are real funcitons

ganeshie8 (ganeshie8):

so, sqrt cannot be negative

OpenStudy (anonymous):

ok